QUE 1a) An investor owns some stock in Harry’s pottery Inc. The stock recently underwent a 5-for -3 stock split . if the stock was trading at $40 per share just before the split, how much is each share most likely selling for after the split? if investor owned 200 shares of stock before the split, how many shares would she own afterward?

   b) Ron’s Rodents Co. has total assets of $5 million, total short and long – term debt of $ 2.8 million and $ 400, 000 worth of 8% preferred stock outstanding. what is the firm’s total book value? What would its book value per share be if the firm had 50, 000 shares of common stock outstanding?
Answer #1

1

how much is each share most likely selling for after the split?

Price after the split=price before the split x(Old shares)(New Shares)

=$40*3/5=$24 per share

if investor owned 200 shares of stock before the split, how many shares would she own afterward?

=200*5/3=333 shares

2

Shareholder Equity= Book value =Total Assets-total short and long – term debt=5-2.8 Million= 2.2 Million or $22,00,000

book value per share=(Shareholder Equity-Less Preferred share)/common stock outstanding

book value per share =(22,00,000- 400, 000)/50000=$36 Per Share